OpenStudy (anonymous):

What is the side length of a cube with a volume of 8000 in3?

OpenStudy (anonymous):

The formula for the volume of a cube is: \[a ^{3}\]

OpenStudy (anonymous):

Volume of cube = side^3

OpenStudy (anonymous):

I SUCK at geometry

OpenStudy (anonymous):

20*20*20=8000

OpenStudy (anonymous):

So side = ---- ?

OpenStudy (anonymous):

thanks guys i dont do very good with this stuff

OpenStudy (anonymous):

So side of cube = 20

OpenStudy (anonymous):

Is it clear to you ?

OpenStudy (anonymous):

Yes yes i gave you a medal :D

OpenStudy (anonymous):

Yes but thats the point, I didnt need the answer i needed help -_-

OpenStudy (anonymous):

No you told it first dude. she told me the formula to help me. she told me after wards because you had already said it.and i didnt ask for the answer.

OpenStudy (anonymous):

And you didnt say 20*20*20=? you said 20*20*20=8000. you gave me the twentys which was the ANSWER -_-

OpenStudy (anonymous):

I cant read the formula you said is all bunched together

OpenStudy (anonymous):

Wow baby

OpenStudy (anonymous):

There ya go!

OpenStudy (anonymous):

my my

OpenStudy (anonymous):

Well, I'm going to delete my stuff, because I really didn't care about the medal.